OpenStudy (anonymous):

d/dx[sqrt(20-x^2-7y^2)]

OpenStudy (anonymous):

1\[-2x^2-14ydy/dx/2\sqrt{?}\]

OpenStudy (anonymous):

sorry i wrote wrong

OpenStudy (phi):

\[ \frac{d}{dx} \sqrt{20-x^2-7y^2} \\ \frac{d}{dx} \left( 20-x^2-7y^2\right)^{\frac{1}{2} } \]

OpenStudy (phi):

start by using the power rule \[ \frac{d}{dx} u^n= n \ u^{n-1} \frac{d}{dx} u \]